A particle with mass 1.72 × 10-3 kg and a charge of 1.17 × 10-8 C has, at a given instant, a velocity v with rightwards arrow on top = (2.77 × 104 m/s)j with hat on top. What is the magnitude of the particle's acceleration produced by a uniform field B with rightwards arrow on top = (1.64 T)i with hat on top +(0.91 T)j with hat on top? (Give your answer in decimal using "m/s2" as unit)
Answer #1

Given
particle mass m = 1.72*10^-3 kg
charge q = 1.17*10^-8 C

velocity vector V = 0 i m/s + 2.77*10^4 j m/s

magnetic field vector B = 1.64 T i + 0.91 T j


the magnetic force on the charge moving inside a magnetic field is

   F = q*V X B

   F = qv*B sin theta

  the particle's acceleration is from Newton's second law F = m*a

a = F/m

a = (5.3150*10^-4)/( 1.72*10^-3) m/s2

a = 0.309011 m/s2